### v2

# An Improved Two-Step Attack on Kyber

Abstract.
After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber was successfully selected in July 2022 and standardized in August 2024. It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase. In this brief, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time. In the first step, we use the correlation power analysis (CPA) to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces. The CPA is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy. In the second step, we adopt the lattice attack to recover s based on the results of CPA. The success rate is largely built up by constructing a trial-and-error method. We implement the proposed attack for the reference implementation of Kyber512 (4 128-value groups of s) on ARM Cortex-M4 and successfully recover a 128-value group of s in about $9$ minutes using a $16$-core machine. Additionally, in that case, we only cost at most $60$ CPA guess values for a group and 15 power traces for a guess.
Lattice-based cryptography, CRYSTALS-Kyber, Side-channel attack, Power analysis
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 2024 ACM/IEEE International Conference on Computer-Aided Design; October 27–31, 2024; NEW JERSEY, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
WITH the fast development of quantum computing, modern public-key cryptographic algorithms are encountering significant threats. Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process in 2016 with the aim of standardizing quantum-resistant cryptographic algorithms. By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and one key encapsulation mechanism (KEM) algorithm after the third round evaluation (alagic2022status, ). CRYSTALS-Kyber is that only KEM, which has been successfully standardized as module-lattice-based KEM (ML-KEM) noted as FIPS 203 from August 2024 (Announcing2024, ). It should be marked that the mathematical security of Kyber has been widely recognized by the cryptography community. However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security (nagata2021physical, ; zhao2023side, ; ravi2024side, ).
Kocher et al. first introduced side-channel attacks (SCAs) in 1996 by leveraging the data dependency on power consumption of cryptographic devices (kocher1996timing, ). Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs, especially to the correlation power analysis (CPA).
Recent years, several CPA attacks on Kyber (ravi2020generic, ; ueno2022curse, ; shen2023find, ; yang2023chosen, ; kuo2023lattice, ) have been proposed by researchers. In 2020, Ravi et al (ravi2020generic, ). used plaintext-checking (PC) oracle-based SCAs to recover the secret key of Kyber512 with 2560 power traces according to the vulnerabilities in Fujisaki-Okamoto (FO) transform for the first time. Then, Ueno et al (ueno2022curse, ). proposed the side channel leakage in re-encryption as a PC oracle, utilizing 3072 power traces to recover the key. The work in (shen2023find, ) deals with imperfect PC oracle SCAs by using the CPA attack twice to recover 60% secret key of Kyber512 at the cost of about 1619 power traces, but there are still 40% error coefficients. Contrast to previous works, in (yang2023chosen, ), Yang et al. tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces. Nevertheless, they had to repeat the attack 256 times to recover the entire secret key of Kyber512. So, they at least need 5120 power traces for a key. In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu presented a two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack. They adopted the CPA attack to recover part of the coefficients of a key and then directly solved the rest coefficients by using a low-dimension equation of M-LWE problem. The correctness was verified by using computer simulations. In this paper, we take the idea of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
Figure 1. The flowchart of the two-step attack.
The proposed two-step attack is summarized in Fig. 1. In the first step, we modify the leakage model to better align our CPA with the arithmetic logic unit. Additionally, we introduce an optional correlation coefficient, the Kendall rank correlation, to address the nonlinearity and robustness issues of the original CPA in real-world attacks. In the second step, we construct a trial-and-error lattice attack algorithm to recover the entire secret key, and optimize the parameters of the algorithm. By employing this two-step attack, we achieve significant improvements in both efficiency and accuracy in practice.
2. Preliminaries
2.1. NTT in CRYSTALS-Kyber
Computing polynomial multiplication in NTT domain can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus significantly improving the speed of the entire encryption and decryption process in Kyber. Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512^{\text{th}}$. Therefore, 256 coefficients are divided into two groups, namely the odd and even groups, for NTT calculations. The even coefficients in NTT domian are described as follows:
(1) $\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\xi^{(2br(i)+1)j}\Rightarrow\ [\hat{f}_{2i}]^{\top}=\mathbf{M}[f_{2i}]^{\top},0\leq i<128,$
where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}$ is a $128\times 128$ matrix with reduced integers. Similarly, INTT can be represented in the same way and the inverse matrix is denoted as $N$.
Listing 1: Base multiplication in Kyber.
1void basemul(int16_t r[2],
2 int16_t zeta,
3 const int16_t s[2],
4 const int16_t u[2])
5{
6 r[0] = fqmul(s[1], u[1]);
7 r[0] = fqmul(r[0], zeta);
8 r[0] += fqmul(s[0], u[0]);
9 r[1] = fqmul(s[0], u[1]);
10 r[1] += fqmul(s[1], u[0]);
11}
2.2. Correlation-Based Attack Techniques
The conventional CPA attack is a non-invasive side-channel attack that uses the Person correlation between power consumption models and actual power traces to recover cryptographic keys. It identifies the correct key by finding the highest correlation between measured and hypothetical values. Pearson correlation coefficient (PCC) is efficient and well-suited for detecting linear relationships between the hypothetical and actual values.
In contrast, Kendall’s tau is a non-parametric measure of association that quantifies the degree of concordance between the rankings of two variables, indicating whether their values are consistently ranked together (abdi2007kendall, ). Kendall’s tau $\tau$ ranges from $-1$ to $1$, with $\tau=1$ indicating perfect agreement, $\tau=-1$ indicating perfect disagreement, and $\tau=0$ indicating no agreement. The Kendall’s tau calculation is computed as follows:
(2) $\tau_{b}=\frac{c-d}{\sqrt{({c+d+t_{x}})({c+d+t_{y}})}},$
where $c$ and $d$ denote the numbers of concordant and discordant pairs, $t_{x}$ and $t_{y}$ represent the counts of tied ranks in the $x$ and $y$ data sets, respectively.
We have separately implemented the two methods in software and found that the PCC computation is much faster than the Kendall’s tau calculation. The main reason is that the latter must compute a large full permutation. Luckily, when we combined the two methods together, the efficiency and reliability both were enhanced. More details will be shown in the following section.
3. Proposed Improved Two-Step Attack
It should be remarked that the proposed improved two-step attack is illustrated in Fig. 1. We will detail the basis of the attack analysis and our proposed three improvements in the following. We assume that an attacker has access to a device which is running Kyber decryption and can only observe ciphertexts entered into the device. In addition, they can also capture the power traces of the device.
Several test vector leakage assessment (TVLA) results from previous works have indicated the potential leakage points during the decryption process in Kyber (shen2023find, ; yang2023chosen, ). One critical side-channel leakage occurs during the point-wise multiplication of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$. In the clean implementation of pqm4 (kannwischer2019pqm4, ), polynomial multiplication employs 256 $basemuls$ in Kyber512. Moreover, each calculation of $basemul$ is independent from the others. From Listing 1, it is clear that the result depends on two secret key coefficients, $s_{0}$ and $s_{1}$. All partial secret key coefficients range from $\left(0,3328\right)$ for Kyber. In a practical attack, suppose that the attacker aims to recover the coefficient $s_{0}$. $r_{0}$ can be selected as the intermediate value.
3.1. Modified Leakage Model for CPA
The Hamming weight (HW) model counts the number of ”1” in intermediate values and is a common linear power model. ARM processors’ ALUs are optimized for two’s complement operations, so we adjust the HW (called modified Hamming weight, MHW) to account for this by using the HW of the 16-bit complement of the intermediate values. This modification improves the accuracy of the HW model and reduces the number of power traces (denoted as $D$) needed for CPA. As shown inside the blue dotted box in Fig. 2 (a) and (b), the correct key coefficient is obscured using the original model (a) but can be effectively identified with just 15 power traces using the modified model (b).
3.2. Optional Kendall’s tau for CPA
When the number of power traces is limited, traditional CPA may face challenges due to high PCC values across guessed keys. To address this, we set a PCC threshold (e.g., 0.9) and extract candidate key coefficients with correlations exceeding this threshold. We then map these candidates to a hypothetical power consumption matrix and compute Kendall’s tau between this matrix and the actual power traces. The candidate with the highest Kendall’s tau value is selected as the correct key coefficient, which helps to reduce false positives. This approach combines PCC for initial filtering with Kendall’s tau for more precise selection, despite Kendall’s tau being more computationally demanding.
Figure 2. (a) and (b) present the PCC values for all guessed keys in a $basemul$ call when $D=15$, for the original and modified models, respectively. (c) shows the ||||PCC|||| values for all guessed keys, while the blue curve in (d) illustrates the occurrence of false positives.
3.3. Trial-and-Error Lattice Attack after CPA
As introduced in (kuo2023lattice, ), suppose that an attacker has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered. Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients. INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ (kuo2023lattice, ), where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients. Similarly, $\mathbf{N}_{B}$ and $\hat{\mathbf{s}}_{iB}$ are obtained in the same manner.
In the above formulas, both $\mathbf{N}_{A}$ and $\hat{\mathbf{s}}_{iA}$ are known. Let $\mathbf{t}=\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$, $\mathbf{A}=-\mathbf{N}_{B}$, and $\mathbf{s}^{\prime}=\hat{\mathbf{s}}_{iB}$. Then, we obtain $\mathbf{t}=\mathbf{A}\cdot\mathbf{s}^{\prime}+\mathbf{s}_{i}\mod q$. This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller. We use Kannan’s embedding technique to solve this problem (kannan1987minkowski, ), where the matrix $\mathbf{B}_{kan}$ is summarized as follows:
[…]
$\left[\mathbf{I}_{sr}\mid\mathbf{A}^{\prime}\right]$ is a reduced row echelon matrix of $\mathbf{A}$ transpose. A vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ of length $129$ is obtained by reduction algorithm, which contains the secret key $\mathbf{s}_{i}^{\top}$ of length 128.
But it is worth noting that the norm of vector $\mathbf{w}$ must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the unique shortest vector problem to be solved. For Kyber512, when $sr\geq 39$, a lattice attack can recover 128 odd/even index coefficients. For Kyber768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered.
We elaborate on the theoretical foundation of lattice attacks, which can tolerate error coefficients of recovered key and still successfully retrieve all 128 key. Leveraging this method, we propose a more flexible and versatile approach for its application. The proposed trial-and-error lattice attack is shown in Listing 2, where $n$ represents the number of CPA, and $\mathbf{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trials, respectively.
One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA instead of $128$ CPA. Certainly, the more CPA are carried out, the higher the success rate of this method will be, but its efficiency will also slow down. The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trial-and-error method. It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trial and lead to the overall recovery inefficient. For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively. Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
Listing 2: Trail-and-Error Lattice Attack.
1index = [1, 2, ..., 37, 38]
2other_index = [39, 40, ..., n-2, n-1]
3guess_sk = [sk_bu0, ..., sk_bu38]
4other_sk = [sk_bu39, ..., sk_bu(n-1)]
5ken = [k_bu0, ..., k_bu38]
6other_ken = [k_bu39, ..., k_bu(n-1)]
7
8result = Lattice_attack(index, guess_sk)
9result = (M @ result)
10if guess_sk == [result[j] for j in index]:
11 return result
12else:
13 for i in count:
14 updata_index(index)
15 updata_sk(guess_sk)
16 result = Lattice_attack(index, guess_sk)
17 result = (M @ result)
18 result_target = [result[j] for j in index]
19 if guess_sk == result_target:
20 return result
21 break
22
23def Lattice_attack(a, b):
24def updata_index():
25def updata_sk():
4. Experiments
The proposed attack is applied to Kyber512, and it is noteworthy that our method is also applicable to other security levels due to their use of the same $basemul$. We run the reference implementation from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz. Power traces were captured using a PicoScope 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe, with a sampling rate set to $500$ MSa/s. Based on the analysis in Section 3, we have chosen $fqmul(k,u_{1})$ as the intermediate value of our attack.
Figure 3. (a) and (b) present the PCC and Ken results, respectively, for a another $basemul$ call when $D=15$. Similarly, (c) and (d) show the PCC and Ken results for the same $basemul$ call when $D=11$.
Step1: CPA
As aforementioned, the CPA is divided into two stages: the capturing stage and the modeling computation stage. Our optimizations are mainly focus on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
For the capturing stage, we carry out the CPA at the NTT operation which calls $basemul$ functions for 128 times. In our CPA experiments, we capture random ciphertexts and tremendous data of electromagnetic radiations during those calls.
To show the effectiveness of the modified leakage model in the modeling computation stage, we have implemented two experiments. The former is shown in Fig. 2 (a) and (b) of Section 3.1 and the modified model can reinforce the PCC value of the correct key. The latter shown in Fig. 2 (c) and (d) is to demonstrate its superiority in identifying the complementary false positive.
When the number of random ciphertexts $D$ is set to $15$, the relationship of the maximum ||||PCC|||| to the guess keys is shown in Fig. 2 (c). We can directly pick out the correct key $2280$ and its complementary false positive $1049$ (equal to $3329-2280$), which is a general phenomenon. In addition, as the number of power traces decreases, the attacker would run into some problems because both the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ would have high and close PCC values, as the HW$(s_{0})$ and HW$(q-s_{0})$ are highly correlated. To escape such false positive, we have modified the leakage model as described in Section 3.1. The correspondence between the PCC values of the correct key coefficient and the false positive one is shown in Fig. 2 (d). It is observed that the positive peak value corresponds to the correct key coefficient and the negative peak value of PCC corresponds to the false positive one. Consequently, to avoid the false positive one is selected, we should focus solely on the positive peak value of PCC. We briefly explain this phenomenon as follows: Let $s_{0}$ be the correct one, and $s_{0}^{\prime}=3329-s_{0}$ be the false positive one. The intermediate values for the two guessed key coefficients are summarized below:
(4) $r_{1}=\text{fqmul}(s_{0},u_{1})=(s_{0}\times u_{1})\,mod\,q,$
(5) $r_{1}^{\prime}=\text{fqmul}(s_{0}^{\prime},u_{1})$ $=((3329-s_{0})\times u_{1})\,mod\,q$
$=-(s_{0}\times u_{1})\,mod\,q.$
MHW$(r_{1})$ and MHW$(r_{1}^{\prime})$ show a negative correlation, where the correct MHW$(r_{1})$ is expected to be a positive PCC. Thus, considering their signs allows us to distinguish them.
To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process. Fig. 3 (a) shows the points of the maximum ||||PCC|||| results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates after PCC is equal to 4. It can be seen that the correlations of four coefficients $564$, $1687$, $2010$, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to $0.930$ and their correlation values are very close. The phenomenon can lead to an inability to distinguish the correct coefficient. This is another kind of false positives. If the attacker just picks one of the four candidates at random, the accuracy is only about $25$%, which is still regarded as a failure. Therefore, we use the method mentioned in Section 3.2 to compensate for such case. First of all, we compute the Kendall coefficient matrix of these four candidate coefficients. It should be noted that the time of this computation is negligible when compared to the complete PCC computing. Fig. 3 (b) shows the values of the four maximum ||||Kendall’s tau|||| when $D=15$. It shows that Kendall’s tau expands the differences between correct coefficient and false positives. As the number of power traces increases, the value of correct coefficient tends to be more stable while the ||||Kendall’s tau|||| of false positives decrease quickly. We can easily distinguish the correct key $564$ when $D=15$ and output it as the final result. In another word, an attacker can take the correct coefficient $564$ for the $basemul$ call using at most 15 power traces. Similarly, we reduce the number of power traces to $11$ step by step and implement the PCC and Kendall’s tau computations. The corresponding results are shown in Fig. 3 (c) and (d). It can be seen that the correct key $564$ can almost not be recognized. We can pick out the final correct candidate from those four candidates according to the results of the summations of their two kinds of corrections. Note that if we continue reducing the number of power traces, we may not be able to figure it out any more.
Figure 4. The success rate and efficiency of our attack in twenty experiments.
Step2: Trial-and-Error Lattice Attack
We conduct the trial-and-error lattice attack using $16$ threads on a Xeon Gold 6342 @ $2.8$ GHz. The parameter of the $block$ size in the BKZ reduction is set to $50$ and the value of $max\_loops$ to $8$. We collect the CPA results of the first sixty $basemuls$ calls. Fig. 4 shows the relationships between the success rate/time and the number of power traces represented in the blue curve and groups of histograms, respectively. The success rate is evaluated by repeatedly conducting the proposed attack many times and counting the success times. It can be seen that when the number of power traces is no less than $15$, all the experimental attacks are successful. Those cases with traces number smaller than $15$ can be compensated to some degree by running more trials of the lattice attack. Meanwhile, according to the decomposed time histograms in Fig. 4, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA is almost unchanged although the traces number is doubled. The main reason could be the different utilization ratios of the multi-threading server. Note that the CPA computations are independently executed with the different power traces while the trial computations of the lattice attack are executed in serial. Therefore, we can summarize that when the success rate is large enough (close/equal to $1$), we can only cost at most $15$ power traces for a guess/call and about $9$ minutes to recover the full key on Kyber512.
Table 1. SCAs on Kyber to Recover a 128-Value Group of $\mathbf{s}$
Work Attack Type No Cip Restriction Attack Target Traces Num for a Call Num of Calls
(xu2021magnifying, ) Pro SPA $\times$ INTT, ME 8$\sim$960 128
(mu2022voltage, ) Pro TA $\times$ MR 11 128
(sim2022chosen, ) Non-Pro SPA $\times$ MR 6$\sim$12 128
(yang2023chosen, ) Non-Pro CPA $\times$ CM 20$\sim$30 128
(kuo2023lattice, ) Non-Pro CPA $\checkmark$ CM 800 50$\sim$87
Ours Non-Pro CPA $\checkmark$ CM 11$\sim$15 60
  • Cip: Ciphertext, Num: Number, Pro: Profile, ME: Message Encoding,
    MR: Modular Reduction, CM: Coefficient Multiplication.
Table 1 shows the comparison results of SCAs on Kyber to recover a 128-value group of $\mathbf{s}$. The attack types include the profiled SPA (xu2021magnifying, ), profiled TA (mu2022voltage, ), non-profiled SPA (sim2022chosen, ), and non-profiled CPA (kuo2023lattice, ; yang2023chosen, ). Our method belongs to the non-profiled CPA, which does not cost extra training step. Moreover, we only adopt random ciphertext, and thus we do not need extra effort like the oracle-based assistance to apply to the KEM scheme (ravi2024side, ). The restriction ciphertext CPA requires oracle-based assistance to be applied to the KEM scheme. For the performance, as shown in the table, the total number of required power traces of ours is $[11,15]\times 60=[660,900]$, much smaller than the state-of-the-art work (yang2023chosen, ) which requires at least $20\times 128=2560$ power traces. If we slightly increase the traces number of the guess, the required time can be reduced to approximate $5$ minutes, much faster than the $20$ minutes recorded in (kuo2023lattice, ).
5. Conclusion
In this brief, we propose an efficient two-step attack including an enhanced CPA and a trial-and-error lattice attack for Kyber. In traditional CPA, false positives are the main factor affecting coefficients of secret key extraction in experiments (yang2023chosen, )and (kuo2023lattice, ). However, by employing the comprehensive CPA step we proposed, we can cleverly avoid false positives while reducing the number of required power traces in the practical attack. In the lattice step, we construct a trial-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time. Experimental results show that our proposed attack can accurately recover the full secret key of Kyber512 in about $9$ minutes with about 15 power traces of a guess on a machine with sixteen threads. Because the $basemul$ function called by different security levels of Kyber is the same, our work can be directly applied to the other parameters. Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography.